# ReorderBench: A Benchmark for Matrix Reordering

### v1

@@ Abstract
Matrix reordering permutes the rows and columns of a matrix to reveal meaningful visual patterns, such as blocks that represent clusters. A comprehensive collection of matrices, along with a scoring method for measuring the quality of visual patterns in these matrices, contributes to building a benchmark. This benchmark is essential for selecting or designing suitable reordering algorithms for specific tasks. In this paper, we build a matrix reordering benchmark, ReorderBench, with the goal of evaluating and improving matrix reordering techniques. This is achieved by generating a large set of representative and diverse matrices and scoring these matrices with a convolution- and entropy-based method. Our benchmark contains 2,835,000 binary matrices and 5,670,000 continuous matrices, each featuring one of four visual patterns: block, off-diagonal block, star, or band. We demonstrate the usefulness of ReorderBench through three main applications in matrix reordering: 1) evaluating different reordering algorithms, 2) creating a unified scoring model to measure the visual patterns in any matrix, and 3) developing a deep learning model for matrix reordering.
@@ I Introduction
Matrix reordering permutes the rows and columns of a matrix to reveal meaningful visual patterns, such as blocks for clusters, off-diagonal blocks for bi-cliques, stars for highly connected nodes, and bands for paths. Despite its significant potential to enhance data visualization and analysis, most existing algorithms are limited by being developed on either a small set of example matrices [1, 2, 3] or large sets of matrices that share the same visual pattern [4]. The lack of extensive and diverse matrices with an appropriate scoring method to measure the quality of their visual patterns makes it difficult to develop an effective reordering method for revealing various patterns. Although Behrisch et al.  [5] have taken the first step in building a matrix reordering dataset, this dataset has two limitations. First, their focus has been limited to binary matrices with entries of 00 or $1$, while real-world applications often involve continuous matrices with real-valued entries. Second, their scoring method is closely tied to the matrix generation process and cannot directly measure the quality of visual patterns in matrices reordered by different algorithms.
@@
To overcome these limitations, it is essential to build a matrix reordering benchmark that includes both binary and continuous matrices, along with a more generic scoring method capable of evaluating the quality of different visual patterns in matrices reordered by different algorithms. Such a benchmark will facilitate the evaluation and development of matrix reordering techniques. First, it enables the selection of suitable reordering algorithms for specific tasks by evaluating their reordering performance. Second, it supports the development of new reordering methods that work well for different visual patterns.
@@
To better support these tasks, we build the ReorderBench benchmark by generating a large collection of representative and diverse symmetric matrices and scoring them with a convolution- and entropy-based method. To ensure the representativeness and diversity of the benchmark, we first generate a set of representative matrix templates for each visual pattern. Then, based on these matrix templates, a large number of matrix variations with diverse degrees of degeneration are generated. The diversity is achieved by combining different variation methods, including noise addition and index swapping. To accurately evaluate the quality of visual patterns in a matrix, we develop a scoring method by combining the matching capability of convolutional kernels [6] and the disorder detection capability of entropy [7]. ReorderBench contains $2,835,000$ binary matrices and $5,670,000$ continuous matrices, categorized into four patterns: block, off-diagonal block, star, and band. Table I provides a statistical comparison with the existing matrix datasets.
@@
TABLE I: A comparison with the existing datasets. ”Quality score?” refers to whether the dataset contains quality scores. ”Direct used?” refers to whether the dataset can be directly used to measure the quality of the reordered matrices.
Dataset # Binary # Continuous Qualtiy score? Direct used?
ReorderBench $2,835,000$ $5,670,000$
@@
  • A matrix reordering benchmark for selecting or designing an appropriate reordering algorithm for specific tasks, available at: https://reorderbench.github.io/.
@@ II Related Work
Our work is related to matrix reordering methods, quality metrics, and matrix dataset.
@@
Matrix reordering method. Matrix reordering is a long-standing research problem that has received attention from multiple fields, such as sociology and bioinformatics [13]. Earlier efforts focus on developing classical algorithms to optimize a given quality metric. According to the optimization strategy, these algorithms can be categorized into two classes: exact algorithms [14, 15] and approximate algorithms [16, 2, 17, 18, 19]. Exact algorithms include the branch-and-bound algorithm [14] and the dynamic programming algorithm [15]. However, because of their exponential time complexity, obtaining exact solutions for matrix reordering becomes infeasible as the matrix size increases. Therefore, approximate algorithms based on various heuristics have been proposed. To place similar rows and columns close to each other, Gruvaeus et al.  [16] reorder matrices through hierarchical clustering. Later studies focus on refining the resulting hierarchical clustering dendrogram using a greedy strategy [2] or simulated annealing [17]. Another line of work utilizes dimension reduction techniques, such as principal component analysis [18] and multi-dimensional scaling [19], to capture the data similarity in a lower-dimensional space. The rows are projected onto one dimension, and the matrix is reordered accordingly. For more details on classical algorithms, the readers are referred to the survey by Behrisch et al.  [20].
@@
Recently, to avoid the trial-and-error process of choosing an appropriate algorithm for the unknown visual patterns in a matrix, several deep learning frameworks have been developed. Building upon dimension-reduction-based reordering methods, Watanabe et al.  [3] use an autoencoder to encode the rows and columns of a given matrix into one-dimensional features. The rows and columns are then reordered in ascending order of the one-dimensional features. Kwon et al.  [4] introduce the use of a variational autoencoder to generate different reordering results for a matrix. However, due to the lack of reordering benchmarks with diverse matrices, existing deep learning models are typically trained to reorder one specific matrix [3, 4]. This limits the generalizability of the developed model. ReorderBench addresses this issue by providing representative and diverse matrices along with accurate quality scores. Upon this benchmark, we build a matrix reordering model that is generalized to previously unseen matrices.
@@
Quality metric. Quality metrics measure the quality of visual patterns in matrices and serve as the optimization criteria for reordering algorithms. Most existing quality metrics implement the idea of placing similar rows and columns close to each other to reveal visual patterns [21]. Generally, they fall into two categories: adjacency-based metrics [22, 23, 24, 25] and dissimilarity-based metrics [26, 27, 28, 29, 2].
Adjacency-based metrics directly take the adjacency matrix as input. Some of these metrics come from the graph drawing area [22]. For instance, the minimum bandwidth (BW) measures the width of the narrowest band needed to enclose all the non-zero entries of the matrix [24]. Other commonly used adjacency-based metrics measure the similarity between adjacent entries in the matrix. An example is the measure of effectiveness (ME) [23], which is calculated as the sum of the product of adjacent entries. In matrices with high measure of effectiveness, the entries are clustered to form block or off-diagonal block patterns.
@@
Dissimilarity-based metrics take as input the dissimilarity matrix, which captures the dissimilarity relationships between rows and columns. Robinson’s seminal work [26] defines a perfectly ordered dissimilarity matrix as an anti-Robinson matrix, where dissimilarity values monotonically increase away from the main diagonal in all rows and columns. Following this, many metrics have been developed to assess deviations from the anti-Robinson matrix, including anti-Robinson events/deviation (AR events/deviation) [27] and gradient measures [28]. Parallel to these metrics, several dissimilarity-based metrics directly measure the distance of the large dissimilarity values from the main diagonal. For example, the linear seriation criterion [29] is defined as the sum of the product of the dissimilarity values and their distances from the main diagonal. Recently, Earle et al.  [2] propose the banded anti-Robinson form (BAR) to better measure the quality of local visual patterns. This metric is a relaxed version of the linear seriation criterion.
@@
These adjacency- or dissimilarity-based metrics work well for block and off-diagonal block patterns. However, they are less effective in measuring other patterns. To ensure that more visual patterns are better revealed, an effective metric is required. To this end, we propose the convolution- and entropy-based scoring method to directly measure the quality of visual patterns in ReorderBench. We also develop a unified scoring model to measure visual patterns in any matrix, including matrices beyond ReorderBench.
@@
Matrix dataset. Many matrix datasets are publicly available, including the Pajek graph collection [8], the Petit Testsuite [9], Matrix Market [10], the SuiteSparse matrix collection [11], and the Network Repository [12]. Existing matrix reordering methods often use some matrices from these datasets to showcase their capability to reorder real-world matrices. However, the lack of quality metrics in these datasets prevents a quantitative evaluation of different reordering methods. Consequently, there is a need for a benchmark that includes both matrices and quality metrics.
@@
Behrisch et al.  [5] have pioneered the creation of the Magnostics dataset, which consists of $5,570$ binary matrices. It aims to evaluate the capability of hand-crafted features in detecting four visual patterns (block, off-diagonal block, star, and band) and two anti-patterns (bandwidth and noise). The dataset is created by first generating a type of visual pattern on empty matrices. The matrices are then gradually degenerated using one of the variation methods, including point-swap, index-swap, and masking. Quality scores are assigned to the matrices based on the level of degeneration introduced by the function, such as the number of index swaps in the index-swap function. This dataset works well to select features capable of detecting visual patterns in matrices. Although the selected features could be used to evaluate reordering algorithms for binary matrices, the authors have not fully explored the exact method. Furthermore, the dataset only includes binary matrices, which limits their usage in real-world applications. In contrast, ReorderBench introduces a convolution- and entropy-based scoring method. Unlike the scoring method used in Magnostics, ours does not rely on the variation methods that degenerate the matrices. As a result, it can be directly used to compare different reordering algorithms. ReorderBench also extends to generating and scoring continuous matrices, addressing the gap left by the Magnostics dataset.
@@ III Visual Pattern Summary
There are four commonly used visual patterns in matrices: block, off-diagonal block, star, and band [20, 21].
@@
Band pattern. A band pattern is characterized by lines that run parallel to the main diagonal of the matrix. It indicates connection paths and transition chains in the underlying graph. For example, in a social network, a path can illustrate the spread of information, opinions, or behaviors through a sequence of links from one node to another.
@@ IV-A Matrix Generation
This step aims to generate a collection of representative and diverse matrices. We achieve representativeness by generating various matrix templates for each visual pattern (template generation). We ensure diversity by degenerating each template into a set of matrix variations through methods such as adding noise and swapping indices (variation generation).
@@ IV-A1 Template Generation
Binary and continuous matrices exhibit similar visual patterns introduced in Sec. III. Thus, the process of generating templates for both types can be largely unified. Previous research on continuous matrices has revealed the distinctiveness of their visual patterns, notably characterized by the Robinson structure [30]. In this structure, entries in a block or star pattern monotonically decrease away from the main diagonal. An off-diagonal block pattern is regarded as a mirrored version of a block pattern, with its larger entries closer to a diagonal that mirrors the main diagonal of the block pattern. For a band pattern, the Robinson structure is not applicable, as its entries are parallel to the main diagonal. This structural property exhibits two advantages in the analysis of matrices. First, it illustrates a key characteristic of fully reordered continuous matrices: similar rows are closely positioned [26]. This aligns with human perception and allows users to better identify visual patterns [25]. Moreover, a matrix featuring this structure is representative of those with non-Robinson structures whose optimal reordering results are the same as this matrix. Second, it facilitates a widely used analytical technique for continuous matrices: filtering low-valued entries to highlight underlying visual patterns [31]. As shown in Fig. 2, the Robinson structure preserves the underlying visual pattern more effectively than the non-Robinson structure when low-valued entries are filtered. A straightforward method for generating continuous templates is to replace the $1$s in a binary template with random values in $[0,1]$. However, this method often destroys the Robinson structures, limiting the representativeness of the resulting templates by excluding those based on Robinson structures. To tackle this issue, we first generate a large set of binary templates and then apply the Robinson structure to produce various continuous ones. This method effectively preserves the representativeness of the templates by including all fully reordered continuous matrices derived from the binary templates.
@@
Continuous templates. To generate a continuous template, we first generate a binary one. Then, as shown in Fig. 4, we apply the Robinson structure and replace the $1$s in a binary template with values in $[0,1]$. The Robinson structure of a block or star pattern is generated by reversing unidimensional scaling, a widely used method in matrix generation [32]. For a pattern with $u$ rows, we first randomly select a set of ascending values $x_{1}\leq x_{2}\leq\dots\leq x_{u}$ from the range $[0,1]$. Then, we replace all $1$s in the pattern, where an entry on the $i$-th row and the $j$-th column is replaced by $1-(x_{i}-x_{j})^{2}$. For an off-diagonal block pattern with $u$ rows and $v$ columns, we first generate a block pattern with $max(u,v)$ rows. We then crop and mirror it to obtain the off-diagonal block pattern. Band patterns are parallel to the main diagonal, preventing the application of the Robinson structure. Thus, the entries in a band pattern are replaced by random values from $[0,1]$.

### v2

@@ Abstract
Matrix reordering permutes the rows and columns of a matrix to reveal meaningful visual patterns, such as blocks that represent clusters. A comprehensive collection of matrices, along with a scoring method for measuring the quality of visual patterns in these matrices, contributes to building a benchmark. This benchmark is essential for selecting or designing suitable reordering algorithms for revealing specific patterns. In this paper, we build a matrix-reordering benchmark, ReorderBench, with the goal of evaluating and improving matrix-reordering techniques. This is achieved by generating a large set of representative and diverse matrices and scoring these matrices with a convolution- and entropy-based method. Our benchmark contains 2,835,000 binary matrices and 5,670,000 continuous matrices, each generated to exhibit one of four visual patterns: block, off-diagonal block, star, or band, along with 450 real-world matrices featuring hybrid visual patterns. We demonstrate the usefulness of ReorderBench through three main applications in matrix reordering: 1) evaluating different reordering algorithms, 2) creating a unified scoring model to measure the visual patterns in any matrix, and 3) developing a deep learning model for matrix reordering.
@@ I Introduction
Matrix reordering permutes the rows and columns of a matrix to reveal meaningful visual patterns, such as blocks for clusters, off-diagonal blocks for bi-cliques, stars for highly connected nodes, and bands for paths. Despite its significant potential to enhance data visualization and analysis, most existing algorithms are limited by being developed for understandable metrics [1, 2, 3] or large sets of matrices that share the same visual pattern [4]. Due to the lack of extensive and diverse matrices with an appropriate scoring method to measure the quality of their visual patterns makes it difficult to develop an effective reordering method for revealing various patterns. Although Behrisch et al.  [5] have taken the first step in building a matrix-reordering dataset, this dataset has two limitations. First, their focus has been limited to binary matrices with entries of 00 or $1$, while real-world applications often involve continuous matrices with real-valued entries. Second, the scoring method in their dataset is based on the level of degeneration introduced during generation, such as the number of index swaps in the index-swap function. This scoring method is intrinsically tied to the matrix-generation process and cannot directly measure the quality of visual patterns in matrices reordered by different algorithms.
@@
To overcome these limitations, it is essential to build a matrix-reordering benchmark that includes both binary and continuous matrices, along with a more generic scoring method capable of evaluating the quality of different visual patterns in matrices reordered by different algorithms. Such a benchmark will facilitate the evaluation and development of matrix-reordering techniques. First, it enables the selection of suitable reordering algorithms for revealing the underlying visual patterns by evaluating their reordering performance. Second, it supports the development of new reordering methods that work well for different visual patterns.
@@
To better support these tasks, we build the ReorderBench benchmark by generating a large collection of representative and diverse symmetric matrices and scoring them with a convolution- and entropy-based method. To ensure the representativeness and diversity of the benchmark, we first generate a set of representative matrix templates for each visual pattern. Then, based on these matrix templates, a large number of matrix variations with diverse degrees of degeneration are generated. The diversity is achieved by combining different variation methods, including noise addition and index swapping. To accurately evaluate the quality of visual patterns in a matrix, we develop a scoring method by combining the matching capability of convolutional kernels [6] and the disorder detection capability of entropy [7]. ReorderBench contains $2,835,000$ binary matrices and $5,670,000$ continuous matrices, each generated to exhibit one of four patterns: block, off-diagonal block, star, or band, along with $450$ real-world matrices featuring hybrid visual patterns. Table I provides a statistical comparison with the existing matrix datasets. A more detailed version of the comparison is provided in Table 1 of the supplemental material.
@@
TABLE I: A comparison with the existing datasets. ”Quality score?” refers to whether the dataset contains quality scores. ”Direct use?” refers to whether its quality score can be directly used to measure the quality of reordered matrices.
Dataset # Binary # Continuous Qualtiy score? Direct use?
ReorderBench $2,835,450$ $5,670,000$
@@
  • A matrix-reordering benchmark for selecting or designing an appropriate reordering algorithm for revealing the underlying visual patterns, available at: https://reorderbench.github.io/.
@@ II Related Work
Our work is related to matrix-reordering methods, quality metrics, and matrix dataset.
@@
Matrix-reordering method. Matrix reordering is a long-standing research problem that has received attention from multiple fields, such as sociology and bioinformatics [13]. Earlier efforts focus on developing classical algorithms to optimize a given quality metric for an individual matrix. According to the optimization strategy, these algorithms can be categorized into two classes: exact algorithms [14, 15] and approximate algorithms [16, 2, 17, 18, 19]. Exact algorithms include the branch-and-bound algorithm [14] and the dynamic programming algorithm [15]. However, because of their exponential time complexity, obtaining exact solutions for matrix reordering becomes infeasible as the matrix size increases. Therefore, approximate algorithms based on various heuristics have been proposed. To place similar rows and columns close to each other, Gruvaeus et al.  [16] reorder matrices through hierarchical clustering. Later studies focus on refining the resulting hierarchical clustering dendrogram using a greedy strategy [2] or simulated annealing [17]. Another line of work utilizes dimension reduction techniques, such as principal component analysis [18] and multi-dimensional scaling [19], to capture the data similarity in a lower-dimensional space. The rows are projected onto one dimension, and the matrix is reordered accordingly. For more details on classical algorithms reordering individual matrices, the readers are referred to the survey by Behrisch et al.  [20]. Another line of work focuses on reordering collections of matrices [21, 22]. For example, Beusekom et al.  [22] aim to achieve contextual orderings that balance the consistency across the collection and the quality of individual reordering results.
@@
Recently, to avoid the trial-and-error process of choosing an appropriate algorithm for the unknown visual patterns in a matrix, several deep learning frameworks have been developed. Building upon dimension-reduction-based reordering methods, Watanabe et al.  [3] use an autoencoder to encode the rows and columns of a given matrix into one-dimensional features. The rows and columns are then reordered in ascending order of the one-dimensional features. Kwon et al.  [4] introduce the use of a variational autoencoder to generate different reordering results for a matrix. However, due to the lack of reordering benchmarks with diverse matrices, existing deep learning models are typically trained on limited datasets that consist of either the given matrix [3] or its reordering results [4]. These training methods aim to learn features specific to the given matrix, resulting in models that are adept at reordering only the given matrix and thus lack generalizability. ReorderBench addresses this issue by providing representative and diverse matrices along with accurate quality scores. Upon this benchmark, we build a matrix-reordering model that is generalized to previously unseen matrices.
@@
Quality metric. Quality metrics measure the quality of visual patterns in matrices and serve as the optimization criteria for reordering algorithms. Most existing quality metrics implement the idea of placing similar rows and columns close to each other to reveal visual patterns [23]. Generally, they fall into two categories: adjacency-based metrics [24, 21] and distance-based metrics [25, 26, 9, 27, 28, 29, 30, 2].
Adjacency-based metrics measure the similarity between adjacent entries in the matrix. An example is the measure of effectiveness (ME) [24], which is calculated as the sum of the product of adjacent entries. In matrices with high measure of effectiveness, the entries are clustered to form block or off-diagonal block patterns. More recently, Moran’s I [21] has been applied to measure the spatial auto-correlation in matrix entries. It classifies adjacencies between entries with values of 0 and 1 into three categories: $1$-$1$, 00-00, and 00-$1$, and calculates a weighted sum of their occurrences. To accommodate matrices with varying levels of sparsity, the weights are determined based on the sum of all entries.
@@
Distance-based metrics focus on the distances between the rows and columns in the ordering. Some of these metrics measure how well the ordering distances align with the connectedness in the underlying graph [25]. For example, the linear arrangement (LA) [9] measures the sum of the distances between the connected vertices in the underlying graph. Other distance-based metrics measure how well the ordering distances align with the dissimilarity between rows and columns based on the dissimilarity matrix. Robinson’s seminal work [27] defines a perfectly ordered dissimilarity matrix as an anti-Robinson matrix, where dissimilarity values monotonically increase away from the main diagonal in all rows and columns. Following this, many metrics have been developed to assess deviations from the anti-Robinson matrix, including anti-Robinson events/deviation (AR events/deviation) [28] and gradient measures [29]. Parallel to these metrics, several distance-based metrics directly measure the distance of the large dissimilarity values from the main diagonal. For example, the linear seriation criterion [30] is defined as the sum of the product of the values and their distance from the main diagonal. Recently, Earle et al.  [2] propose the banded anti-Robinson form (BAR) to better measure the quality of local visual patterns. This metric is a relaxed version of the linear seriation criterion.
@@
These adjacency- or distance-based metrics work well for block and off-diagonal block patterns. However, as they typically focus on placing similar rows and columns close to each other, they are less effective in measuring other patterns, such as star patterns and band patterns, where even small amounts of noise can notably distort row and column similarities. To ensure that more visual patterns are better revealed, an effective metric is required. To this end, we propose the convolution- and entropy-based scoring method, which combines the matching capability of convolutional kernels [6] and the disorder detection capability of entropy [7] to directly measure the quality of visual patterns in ReorderBench. This scoring method accurately evaluates the quality of visual patterns in a matrix, and is essential for constructing a benchmark that supports the evaluation and development of matrix-reordering techniques. We also develop a unified scoring model to measure visual patterns in any matrix, including matrices beyond ReorderBench.
@@
Matrix dataset. Many matrix datasets are publicly available, including the Pajek graph collection [8], the Petit Testsuite [9], Matrix Market [10], the SuiteSparse matrix collection [11], and the Network Repository [12]. Existing matrix-reordering methods often use some matrices from these datasets to showcase their reordering capability. However, the lack of quality metrics in these datasets prevents a quantitative evaluation of reordering methods. Consequently, there is a need for a benchmark that includes both matrices and quality metrics.
@@
Behrisch et al.  [5] have pioneered the creation of the Magnostics dataset, which consists of $5,570$ binary matrices. It aims to evaluate the capability of hand-crafted features in detecting four visual patterns (block, off-diagonal block, star, and band) and two anti-patterns (bandwidth and noise). The dataset is created by first generating a type of visual pattern on empty matrices. The matrices are then gradually degenerated using one of the variation methods, including point-swap, index-swap, and masking. Quality scores are assigned to the matrices based on the level of degeneration introduced by the function, such as the number of index swaps in the index-swap function. This dataset works well to select high-dimensional features capable of detecting visual patterns in matrices using distance-based search. Although the selected features could potentially be used to derive a quality metric for evaluating reordering algorithms for binary matrices, the authors have not fully explored the exact method. Furthermore, the dataset only includes binary matrices, which limits their usage in real-world applications. In contrast, ReorderBench introduces a convolution- and entropy-based scoring method. Unlike the scoring method used in Magnostics, it does not rely on the variation methods that degenerate the matrices. As a result, it can be directly used to compare reordering algorithms. ReorderBench also extends to generating and scoring continuous matrices, addressing the gap left by the Magnostics dataset.
@@ III Visual Pattern Summary
There are four commonly used visual patterns in matrices: block, off-diagonal block, star, and band [20, 23, 5, 21]. Despite the ongoing debate regarding the interpretation of band [31], we include it in our work due to its application in several important fields, such as bioinformatics [32] and network analysis [33]. This inclusion also serves to demonstrate the generalizability of our matrix generation and scoring methods. For applications where the band patterns are not required, they can be excluded from our benchmark.
@@
Band pattern. A band pattern is characterized by lines that run parallel to the main diagonal of the matrix. It indicates the presence of paths, cycles, or meshes in the underlying graph. For example, in a social network, a path can illustrate the spread of information, opinions, or behaviors through a sequence of links from one node to another.
@@ IV-A Matrix Generation
This step aims to generate a collection of representative and diverse matrices with non-overlapping patterns. Since there is a one-to-one correspondence between these matrices and their underlying graphs, this process also ensures the representativeness and diversity of graph structures. We achieve representativeness by generating various matrix templates for each visual pattern (template generation). We ensure diversity by degenerating each template into a set of matrix variations through methods such as adding noise and swapping indices (variation generation).
@@ IV-A1 Template Generation
Binary and continuous matrices exhibit similar visual patterns introduced in Sec. III. Thus, the process of generating templates for both types can be largely unified. Previous research on continuous matrices has revealed the distinctiveness of their visual patterns, notably characterized by the Robinson structure [34]. In this structure, entries in a block or star pattern monotonically decrease away from the main diagonal. An off-diagonal block pattern is regarded as a mirrored version of a block pattern, with its larger entries closer to a diagonal that mirrors the main diagonal of the block pattern. For a band pattern, the Robinson structure is not applicable, as its entries are parallel to the main diagonal. This structural property exhibits two advantages in the analysis of matrices. First, it illustrates a key characteristic of fully reordered continuous matrices: similar rows are closely positioned [27]. This aligns with human perception and allows users to better identify visual patterns [21]. Moreover, a matrix featuring this structure is representative of those with non-Robinson structures whose optimal reordering results are the same as this matrix. Second, it facilitates a widely used analytical technique for continuous matrices: filtering low-valued entries to highlight underlying visual patterns [35]. As shown in Fig. 2, the Robinson structure preserves the underlying visual pattern more effectively than the non-Robinson structure when low-valued entries are filtered. A straightforward method for generating continuous templates is to replace the $1$s in a binary template with random values in $[0,1]$. However, this method often destroys the Robinson structures, limiting the representativeness of the resulting templates by excluding those based on Robinson structures. Previous works have shown that starting with representative examples and diversifying them is a practical method for constructing effective datasets [36, 37]. Thus, to obtain a set of representative continuous matrix templates, we first generate a large set of binary templates and then apply the Robinson structure to produce various continuous ones. This method effectively preserves the representativeness of the templates by including all fully reordered continuous matrices derived from the binary templates.
@@
Continuous templates. To generate a continuous template, we first generate a binary one. Then, as shown in Fig. 4, we apply the Robinson structure and replace the $1$s in a binary template with values in $[0,1]$. The Robinson structure of a block or star pattern is generated by reversing unidimensional scaling, a widely used method in matrix generation [39]. For a pattern with $u$ rows, we first randomly select a set of ascending values $x_{1}\leq x_{2}\leq\dots\leq x_{u}$ from the range $[0,1]$. Then, we replace all $1$s in the pattern, where an entry on the $i$-th row and the $j$-th column is replaced by $1-(x_{i}-x_{j})^{2}$. For an off-diagonal block pattern with $u$ rows and $v$ columns, we first generate a block pattern with $max(u,v)$ rows. We then crop and mirror it to obtain the off-diagonal block pattern. Band patterns are parallel to the main diagonal, preventing the application of the Robinson structure. Thus, the entries in a band pattern are replaced by random values from $[0,1]$. For applications where values above a certain threshold are considered meaningful, such as those interpreting entries as strength or likelihood of connectivity, the entries can be sampled from a constrained range (e.g., [0.1,1]) to strictly ensure that they exceed a minimum threshold.